One of the legs of a right triangle is twice as long as the other, and the perimeter is 28. Find the lengths of all 3 sides, to three decimal places.
Answer by richwmiller(17219)

x^2+2x^2=c^2
c=sqrt(x^2+(2x)^2)
c=sqrt(x^2+4x^2)
c=sqrt(5x^2)
c=sqrt(5)x
x+2x+sqrt(5)x=28,
y=2x
z=sqrt(5)x
x=5.3475
x = -7 (sqrt(5)-3), y = -14 (sqrt(5)-3), z = 7 (3 sqrt(5)-5)
x=5.3475, y=10.695, z=11.957
x^2+y^2=z^2